OpenStudy (anonymous):

100 members in a spot club 57 play cricket 50 '' football 48 '' hockey 19 '' cricket and football 23 '' football and hockey 21 '' play cricket and hockey find the valve of x find members of members playing foot ball only

OpenStudy (anonymous):

Where is x ?

OpenStudy (anonymous):

i dont kwn where is x that what got 4 home work just cant do it

OpenStudy (sumeer):

 data we have a+b+c+d+e+f+g=100 a+b+c+d=50 c+d+e+g=48 b+d+e+f=57 b=19 c=23 e=21 solution a+d+f+g+(b+c+e)=37+(63) d+f=17 look b+d+e+f a+d=18 from a+b+c+d a+g=20 from a+d+f+g a+2d+g=22 from a+b+c+d and c+d+e+g 2d=2 and d=1 a+d=18 => a=17

OpenStudy (anonymous):

 | the veen diagram above illustrates some of the information given below there are 100 members in a foreign language club 49 members speak spanish 19 '' '' spanish and arabic 53 '' '' arabic 45 '' '' french 15 '' '' spanish and french 22 '' '' french '' arabic each member speaks at least one of the three languages a) lat the number of members who speak all three languages be x 1) write an algebraic expression to represent the number of students in the shaded region a) deseribe the region shaded b) write an equation to show the total number of members in the club c)determine the member of members who speak all three language